Raner, Harris & Chan is a consulting firm that specializes in information systems for medical and dental clinics. The firm has two offices—one in Chicago and one in Minneapolis. The firm classifies the direct costs of consulting jobs as variable costs. A contribution format segmented income statement for the company’s most recent year is given:

Office
Total Company Chicago Minneapolis
Sales $ 712,500 100.0 % $ 142,500 100 % $ 570,000 100 %
Variable expenses 384,750 54.0 % 42,750 30 % 342,000 60 %
Contribution margin 327,750 46.0 % 99,750 70 % 228,000 40 %
Traceable fixed expenses 159,600 22.4 % 74,100 52 % 85,500 15 %
Office segment margin 168,150 23.6 % $ 25,650 18 % $ 142,500 25 %
Common fixed expenses not traceable to offices 114,000 16.0 %
Net operating income $ 54,150 7.6 %

Required:

1-a. Compute the companywide break-even point in dollar sales.

1-b. Compute the break-even point for the Chicago office and for the Minneapolis office.

1-c. Is the companywide break-even point greater than, less than, or equal to the sum of the Chicago and Minneapolis break-even points?
Answer #1

1a.

Breakeven point in dollars = ($159,600 + $114,000) / 46% = $594,783

1b.

Breakeven point:

Chicago = $74,100 / 70% = $105,857

Minneapolis = $85,500 / 40% = $213,750

1c.

Companywide Breakeven point = $594,783

Total of Chicago and Minneapolis Breakeven point = $105,857 + $213,750 = $319,607

Companywide Breakeven point is greater.